# The Magnetic Filling in MAD Simulations and its Impact on the Jet in M87

### v1

@@ 2.1 Resolution and MRI Suppression Tests
In this section we investigate the resolution of the magneto-rotational instability (MRI), which is thought to be the main driver of angular momentum and energy transport of plasma accretion in SMBHs systems. Due to the limitations of finite resolution, the spectrum of MRI modes is truncated in simulations, hence is able to produce various numerical artifacts. Truncation of the MRI modes affects the wavenumber range of turbulence modes caused by the MRI, and consequently alters the rate at which nonlinear mode-mode couplings transfer energy from large- to small-scale dynamics. Insufficient resolution of the fastest growing MRI mode brings the turbulent field amplification to a halt and numerical dissipation takes over, hence the field decays. We test wether underlying MRI modes are well resolved by the physical criterion, that the number of grid-cells that are able to capture the MRI wavelength $\lambda_{\mathrm{MRI}}$ is sufficiently large. The so-called quality- or $\mathcal{Q}$-factors (see Sano et al. (2004); Noble et al. (2010); Porth et al. (2019)) are defined to reflect the resolution of the underlying MRI mode with the grid spacing in a given spatial direction $\Delta x^{(i)}$:
@@
The wavelength of the fastest growing MRI mode and the cell extent in a given direction can be defined as:
@@
$\Omega=u^{\phi}/u^{t}$ is the azimuthal angular velocity of the fluid measured in the locally non-rotating reference frame (LNRF) in Kerr-spacetime and $e_{\mu}^{(i)}$ is the tetrad basis one-form, where the LNRF is boosted into the comoving frame of the fluid (cf. Takahashi 2008, for the basis transformations). $\Delta x^{(i)}$ is the grid spacing in a fluid frame tetrad basis oriented along the LNRF. $(\Delta x^{\mu})^{(i)}$ is a vector containing containing for each $i$ one of the grid spacings in Kerr-Schild coordinates. For the $\theta$-coordinate and grid spacing $\Delta\theta$ in Kerr-Schild this would look like the following: $(\Delta x^{\mu})^{(\theta)}=(0,0,\Delta\theta,0)^{T}$ and is analogously calculated for the other coordinates $(r,\phi)$. We calculate the Quality factors $Q^{r}$, $Q^{\theta}$ and $Q^{\phi}$ and average them over a region that includes the wind and disk regions, which is defined by the bounds $\pi/3\geq\theta_{\mathrm{disk}}\geq 2\pi/3$ and radial upper bound $r_{\mathrm{disk}}\leq 50\,\mathrm{M}$, where we always take the full azimuthal map of $0\geq\phi>2\pi$. The spatial average is calculated as the expectation value over the covariant volume as follows:
@@
where $w(r,\theta,\phi)$ is a weighting function that acts as a masking function. In the case of the disk averages this function is 1 inside the poloidal bounds specified above and 0 outside.
We show the temporal evolution of the disk averged quality-factors and their median values in Appendix D Fig. 18. Sano et al. (2004) suggested $\mathcal{Q}$-factors of $\mathcal{Q}>6$ to corresponde to sufficient resolution of the MRI. We match these requirements in $Q^{(\theta)}$ and $Q^{(\phi)}$ with consistent values in all simulations with median values in time of $Q_{\mathrm{med}}^{(\theta)}\gtrsim 35,Q_{\mathrm{med}}^{(\phi)}\gtrsim 12$ and time averages of $\langle Q^{(\theta)}\rangle_{t}\gtrsim 60,\langle Q^{(\phi)}\rangle_{t}\gtrsim 16$ valid for all simulations over the time span ranging from [10000,30000][M][10000,30000]\,[M] (exact values are shown in Fig. 18 and meridional snapshots in Appendix D Fig. 19).
@@ 2.2 Electromagnetic and Kinetic Energy Flux in the Jet
Figure 3: Temporal evolution of the total energy flux (solid lines) and EM energy flux (dashed lines) measured at $r=10\,\mathrm{M}$ in the jet for simulations K.60 (orange), K.80 (green), K.90 (blue), K.95 (magenta) and K.99 (red). The horizontal lines show the average of the total energy fluxes for each run over the shown time interval.
Figure 4: Time averaged total energy fluxes (dash-dotted lines), EM energy fluxes (solid lines) and kinetic energy fluxes (dashed lines) in the time interval (2500030000)M(25000-30000)\,M for varying BH separations along the jet for simulations K.60 (orange), K.80 (green), K.90 (blue), K.95 (magenta) and K.99 (red). The vertical dotted lines indicate the radii at which the kinetic and EM energy fluxes are in equipartition.
For the time-series analysis, presented henceforth, we estimate the error of all time-averaged quantities by means of the correlation error $\sigma_{\rm{corr}}$ (cf. Appendix E for the formular and more details). This takes into account that the number of data points that are independent of one another is reduced if the data is correlated. The total energy flux is determined via the energy momentum tensor from Eq. (2) by contraction with the metric tensor and reformulating the enthalpy in terms of the internal energy density $u=\nicefrac{{p}}{{(\hat{\gamma}-1)}}$ to be
$T^{\nu}_{\mu}$ $=(\rho+\hat{\gamma}u+b^{2})\,u^{\nu}u_{\mu}\hskip 12.0pt-b^{\nu}b_{\mu}+(p+\nicefrac{{b^{2}}}{{2}})\delta^{\nu}_{\mu}$ (11)
@@
where $\delta^{\nu}_{\mu}$ is the Kronecker delta. The sub tensor responsible for the EM contribution to the $r$-$t$ component is
@@
And analogously the sub tensor for the kinetic contributions is
@@
The total, EM and kinetic energy fluxes are then determined by
$\dot{E}$ $=-\int\hskip-6.0pt\int\sqrt{-g}\mathrm{d}\theta\mathrm{d}\phi\,T_{t}^{r}$ (15)
$\dot{E}_{\mathrm{EM}}$ $=-\int\hskip-6.0pt\int\sqrt{-g}\mathrm{d}\theta\mathrm{d}\phi\,T_{\mathrm{EM}\,t}^{\hskip 11.0ptr}$ (16)
$\dot{E}_{\mathrm{Kin}}$ $=-\int\hskip-6.0pt\int\sqrt{-g}\mathrm{d}\theta\mathrm{d}\phi\,T_{\mathrm{Kin}\,t}^{\hskip 11.0ptr}$ (17)
Before executing the integral we impose the condition that the Bernoulli parameter has to be >1.02>1.02, corresponding to the outflowing part of the plasma shown by the area inside the contours in the bottom row of Fig. 10. We further calculate the energy fluxes at varying radial coordinates and times.
In Fig. 3 we show the time dependence of the total and EM energy flux calculated at a fixed radius $r=10\,\mathrm{M}$. K.99 on average exhibits a significantly higher energy flux of $\sim 50\,[\mathrm{a.u.}]$ exceeding the other simulations by $\gtrsim 3\sigma_{\rm{corr}}$, directly boosting the efficiency of the jet launching, due to the comparable mass accretion rates over all simulations. The energy flux at the jet base is highly dominated by the EM part of the energy flux ($\sim 87\,\%$) for K.99 and causes the total energy flux excess. The dependence of the time-averaged total energy flux through the jet on the distance from the BH along the jet parallel axis can be taken from Fig. 4. The EM and kinetic contributions are in equipartition at different distances from the BH (vertical dashed lines) for each simulation. This is especially evident for simulation K.99, being significantly more electromagnetically dominated both in extent along the jet and magnitude. We note, that the energy fluxes increase with BH seperation along the jet, since plasma from the wind region becomes graviationally unbound, i.e. it reaches $hu_{t}>1.02$.
@@ 2.3 Outflow Efficiencies
Figure 5: Jet efficiencies for simulations K.60 (orange), K.80 (green), K.90 (blue), K.95 (magenta) and K.99 (red), where the dashed horizontal lines show the time averaged efficiencies and the correlation corrected $2\sigma_{\rm{corr}}$ error-band over the shown interval (listed in Tab. 1).
@@
where we make use of Eq. (15) and calculate the mass accretion rate at $r=5\,\mathrm{M}$, $\dot{M}_{\mathrm{5M}}$, and its average $\langle\dot{M}_{\mathrm{5M}}\rangle$ over the time interval $t=[25000,\,30000]\,\mathrm{M}$. This time-intervall is best suited, as the accretion disk is well converged into a steady-state. This can be seen by the mass-accretion rate deccellerating its decline towards the end of the simulations (cf. Fig. 2). For comparison, we also show $\dot{M}_{\rm{5M}}$ in the beforementioned interval in Appendix Fig. 17. The nearly identical average mass-accretion rates as well aid the comparison of the efficiencies (see Eq. (18)).
We explicitly do not use the mass accretion rate at the event horizon $\dot{M}$ (compare Fig. 2), but instead calculate the flux at $5\,\mathrm{M}$ since this excludes the injected matter to set the density floor at the inner simulation boundary. The radial rest-mass energy density flux is defined as follows:
$\dot{M}_{\mathrm{5M}}=\int\hskip-6.0pt\int_{r=\mathrm{5M}}\sqrt{-g}\mathrm{d}\theta\mathrm{d}\phi\,\rho u^{r}$ (19)
The resulting efficiencies of the energy extraction process in all three models are shown in Fig. 5. On average the K.99 simulation exhibits a vastly improved efficiency of $\sim 60\,\%$ excess compared to the commonly used MAD-setup K.80. The averages, errors and peak values of all simulations can be found in Tab. 1.
@@
Table 1: Time-averages, corresponding error and maxima of efficiencies in the time interval $t=\{25000\,\mathrm{M}-30000\,\mathrm{M}\}$ for each GRMHD simulation (cf. Fig. 5).
Figure 6: Left panel: Time and azimuthally averaged density distribution in the meridional plane and total angular momentum flux streamlines (color-coded by a log-scale of the absolute value; increasing from black to white). Middle column: Polar slices of the radial angular momentum flux at $r=2\,\mathrm{M}$, indicated by the dashed, grey lines in the left panel and right column. Shown are total, advective and Maxwell-stress-induced angular momentum fluxes from top to bottom. Right column: Shell-integrated, radial component of the angular momentum fluxes. Solid lines indicate positive values, i.e. outward fluxes and dashed lines are the absolute values of inward (negative) flux.
2.4 Angular Momentum Transport
In the following section, we investigate the angular momentum transport, by that linking it to the jet efficiency as well as to the impact of the magentic energy content in the torus on the dynamics from the initialization of the simulations. Following the approach and definitions of the angular momentum fluxes of Chatterjee & Narayan (2022), we define the radial component of the time averaged and shell-integrated angular momentum flux vectors $\dot{J}^{i}$ ($i\in(r,\theta)$), accordingly. The total flux is defined as
@@ 2.4 Angular Momentum Transport
the advective flux reads
@@
The stress induced flux is given by
@@
which can be further subdivided into the flux induced by Maxwell-stresses
@@
and by Reynolds-stresses, which are subdominant and therefore do not add to the discussions in this paper, hence $\dot{J}^{i}_{\rm{stress}}\sim\dot{J}^{i}_{\rm{stress,M}}$.
@@
We time-average the quantities again over the interval $t=[25000,\,30000]\,\rm{M}$ in order to compare it with the jet energetics and efficiencies from the previous Sec. 2.2 and Sec. 2.3.
@@
In Fig. 6 we present the polar and radial dependence of time- and azimuthally averaged angular momentum transport. In the left panel we show the time- and azimuthally averaged total angular momentum vector-field for the first time for simulations in a MAD-setup with a high prograde BH spin (cf. Chatterjee & Narayan (2022) Fig. 10, showing the case of a non-rotating BH in the MAD-setup). The meridional cross-section of the angular momemtum vector-field exhibits quite distinct characteristica from the case of the MAD-setup with a non-rotating BH, in that, it exhibits a monopol field-geometry in the jet and wind-region up to a polar angle of $\sim 70^{\circ}$ and transitioning into a quadrupol field-geometry in the disk and torus. Through this magnetospheric configuration, the polar outflow from jet and wind feeds back into the equatorial inflow in the, opposed to the non-rotating MADs, where there is a transition from an equatorially parallel inflow inside the disk scale height (similar opening angle as the transition to the loops) to the parabolically outflowing angular momentum flux in the jet. The major difference here is the dominance of the angular momentum transport driven by the rotating BH, where the dominantly radial outflow is redirected towards the equatorial plane into the disk, where the torus has to facilitate the radially-outward angular momentum transport in the wind region.
@@
In the middle and right column of Fig. 6, we investigate the impact of the magnetic filling of the MAD-torus on the angular momentum transport. The middle column shows the polar dependence of the time- and azimuthally averaged, radial component of the angular momentum vectors at $r=2$ M, where there is a maximum in the shell-integrated fluxes. The right panel shows the shell-integrated (see Sec. 2.2 for the definition of shell-integration) and time averaged, radial angular momentum fluxes. The polar slices (indicated by grey dashed lines in the left panel and right column) of $\dot{J}_{\rm{tot}}^{r}$ and $\dot{J}_{\rm{stress,M}}^{r}$ show a clear model trend. Higher initial magnetic energy from the torus leads to an increase outward angular momentum flux, which is also sustained to larger radii $r\lesssim 50\,\mathrm{M}$, presented in Appendix Fig. 20.
@@
In the right panel at the top, one can see that the total shell-integrated fluxes are for one, positive, i.e. radially outgoing flux in the jet and wind are dominant, whereas for the case of a non-rotating BH in Chatterjee & Narayan (2022) Fig. 14, the shell-integrated radial component of the total angular momentum flux is constant and negative up to $r\sim 100$ M. Secondly, the angular momentum fluxes overall peak close to the event horizon of the BH, specifically, inside the ergosphere and drops outside of this region.
@@
We note, that the model-trend is not identical for the shell-integrated total and Maxwell-induced fluxes, since these are affected by the inward advective flux, that is more sensitive to turbulent realisations and does not appear to exhibit any model trend. The model trend is the clearest in the jet sheath opposed to the equatorial plane in both total and Maxwell-induced angular momentum fluxes. Looking at the middle panel, one can clearly see, that the advective, equatorial flux for K.80 is the highest, hence decreasing the total, shell-integrated flux the most for this model. However, we observe that simulation K.99 overall exhibits the highest angular momentum transport both inward, advectively- and predominantly outward, electro-magnetically-driven in jet- and wind-regions. Further we observe, that the model-trend in the jet efficiency and the shell-integrated radial angular momentum fluxes is consistent with one another, which directly links the efficiencies to the angular momentum transport of these simulations.

### v2

@@ 2.1 Resolution and MRI suppression tests
In this section, we describe our investigation of the MRI resolution, which is thought to be the main driver of angular momentum and energy transport of plasma accretion in SMBH systems. Due to the limitations of finite resolution, the spectrum of MRI modes is truncated in simulations; hence, it is able to produce various numerical artifacts. Truncation of the MRI modes affects the wave-number range of turbulence modes caused by the MRI and consequently alters the rate at which nonlinear mode–mode couplings transfer energy from large- to small-scale dynamics. Insufficient resolution of the fastest growing MRI mode brings the turbulent field amplification to a halt, and numerical dissipation takes over; hence, the field decays. We tested wether underlying MRI modes were well resolved by the physical criterion and that the number of grid cells that are able to capture the MRI wavelength $\lambda_{\mathrm{MRI}}$ was sufficiently large. The so-called quality or $Q$ factors (Sano et al. (2004); Noble et al. (2010); Porth et al. (2019)) were defined to reflect the resolution of the underlying MRI mode with the grid spacing in a given spatial direction, $\Delta x^{(i)}$:
@@
The wavelength of the fastest growing MRI mode and the cell extent in a given direction can be defined as
@@
$\Omega=u^{\phi}/u^{t}$ is the azimuthal angular velocity of the fluid measured in the locally non-rotating reference frame (LNRF) in Kerr space time, and $e_{\mu}^{(i)}$ is the $i$-th one-form of the co-vielbein, determining the tetrad basis, where the LNRF is boosted into the comoving frame of the fluid (see Takahashi (2008) for the basis transformations). $\Delta x^{(i)}$ is the grid spacing in a fluid frame tetrad basis oriented along the LNRF. $(\Delta x^{\mu})^{(i)}$ is a vector containing, for each $i,$ one of the grid spacings in the Kerr–Schild coordinates. For the $\theta$ coordinate and grid spacing, $\Delta\theta,$ in Kerr–Schild this would look as follows: $(\Delta x^{\mu})^{(\theta)}=(0,0,\Delta\theta,0)^{T}$. This is analogously calculated for the other coordinates $(r,\phi)$. We calculated the quality factors $Q^{\theta}$ and $Q^{\phi}$, and we averaged them over a region that includes the wind and disk regions, which is defined by the bounds $\pi/3\geq\theta_{\mathrm{disk}}\geq 2\pi/3$ and the radial upper bound $r_{\mathrm{disk}}\leq 50\,\mathrm{M}$, where we always took the full azimuthal map of $0\geq\phi>2\pi$. The spatial average is calculated as the expectation value over the covariant volume as follows:
@@
where $w(r,\theta,\phi)$ is a weighting function that acts as a masking function. In the case of the disk averages, this function is 1 inside the poloidal bounds specified above and 0 outside of them.
We show the temporal evolution of the disk’s averaged quality-factors and their median values in Appendix D, Fig. 18. Sano et al. (2004) suggested $Q$ factors of $Q>6$ to correspond to a sufficient resolution of the MRI. We matched these requirements in $Q^{(\theta)}$ and $Q^{(\phi)}$ with consistent values in all simulations with median values in time of $Q_{\mathrm{med}}^{(\theta)}\gtrsim 35,Q_{\mathrm{med}}^{(\phi)}\gtrsim 12$ and time averages of $\langle Q^{(\theta)}\rangle_{t}\gtrsim 60,\langle Q^{(\phi)}\rangle_{t}\gtrsim 16,$ which are valid for all simulations over the time span $10000,30000]\,\rm M$ (exact values are shown in Fig. 18, and meridional snapshots are given in Fig. 19).
@@ 2.2 Electromagnetic and kinetic energy flux in the jet
Figure 3: Temporal evolution of total energy flux (solid lines) and EM energy flux (dashed lines) measured at $r=10\,\mathrm{M}$ in the jet for simulations K.60 (orange), K.80 (green), K.90 (blue), K.95 (magenta), and K.99 (red). The horizontal lines show the average of the total energy fluxes for each run over the shown time interval.
For the time-series analysis presented here, we estimated the error of all time-averaged quantities by means of the correlation error, $\sigma_{\rm{corr}}$ (see Appendix E for the formula and more details). This takes into account that the number of data points that are independent of each other is reduced if the data are correlated. The total energy flux is determined via the energy momentum tensor from Eq. (2) by contraction with the metric tensor and reformulating the enthalpy in terms of the internal energy density $u=\nicefrac{{p}}{{(\hat{\gamma}-1)}}$ to be
$T^{\nu}_{\mu}$ $=(\rho+\hat{\gamma}u+b^{2})\,u^{\nu}u_{\mu}\hskip 12.0pt-b^{\nu}b_{\mu}+(p+\nicefrac{{b^{2}}}{{2}})\delta^{\nu}_{\mu},$ (11)
@@
where $\delta^{\nu}_{\mu}$ is the Kronecker delta. The sub-tensor responsible for the EM contribution to the $r$-$t$ component is
@@
Analogously, the sub-tensor for the kinetic contributions is
@@
The total, electromagnetic (EM), and kinetic energy fluxes were then determined by
$\dot{E}$ $=-\int\hskip-6.0pt\int\sqrt{-g}\mathrm{d}\theta\mathrm{d}\phi\,T_{t}^{r},$ (15)
$\dot{E}_{\mathrm{EM}}$ $=-\int\hskip-6.0pt\int\sqrt{-g}\mathrm{d}\theta\mathrm{d}\phi\,T_{\mathrm{EM}\,t}^{\hskip 11.0ptr},$ (16)
$\dot{E}_{\mathrm{Kin}}$ $=-\int\hskip-6.0pt\int\sqrt{-g}\mathrm{d}\theta\mathrm{d}\phi\,T_{\mathrm{Kin}\,t.}^{\hskip 11.0ptr}$ (17)
Before executing the integral, we imposed the condition that the Bernoulli parameter had to be >1.02>1.02, corresponding to the outflowing part of the plasma shown by the area inside the contours in the bottom row of Fig. 10. We further calculate the energy fluxes at varying radial coordinates and times.
In Fig. 3, we show that the time dependence of the total and EM energy fluxes calculated at a fixed radius of $r=10\,\mathrm{M}$. K.99 on average exhibits a significantly higher energy flux of $\sim 50\,[\mathrm{a.u.}],$ exceeding the other simulations by $\gtrsim 3\sigma_{\rm{corr}}$; this directly boosts the efficiency of the jet launching due to the comparable mass-accretion rates over all simulations. The energy flux at the jet base is highly dominated by the EM part of the energy flux ($\sim 87\,\%$) for K.99 and causes the total energy flux excess. The dependence of the time-averaged total energy flux through the jet on the distance from the BH along the jet’s parallel axis can be taken from Fig. 4. The EM and kinetic contributions are in equipartition at different distances from the BH (vertical dashed lines) for each simulation. This is especially evident for simulation K.99, which is significantly more electromagnetically dominated both in extent along the jet and magnitude. We note that the energy fluxes increase with BH separation along the jet, since plasma from the wind region becomes gravitationally unbound; i.e., it reaches $hu_{t}>1.02$.
Figure 4: Time-averaged total energy fluxes (dash-dotted lines), EM energy fluxes (solid lines), and kinetic energy fluxes (dashed lines) in the time interval (2500030000)M(25000-30000)\,M for varying BH separations along the jet for simulations K.60 (orange), K.80 (green), K.90 (blue), K.95 (magenta), and K.99 (red). The vertical dotted lines indicate the radii at which the kinetic and EM energy fluxes are in equipartition.
@@ 2.3 Outflow efficiencies
Figure 5: Jet efficiencies for simulations K.60 (orange), K.80 (green), K.90 (blue), K.95 (magenta), and K.99 (red), where the dashed horizontal lines show the time-averaged efficiencies and the correlation-corrected $2\sigma_{\rm{corr}}$ error band over the shown interval (listed in Tab. 1).
@@
where we made use of Eq. (15) and calculated the mass-accretion rate at $r=5\,\mathrm{M}$, $\dot{M}_{\mathrm{5M}}$ and its averaged $\langle\dot{M}_{\mathrm{5M}}\rangle$ over the time interval $t=[25000,\,30000]\,\mathrm{M}$. This time interval is the most suitable one, as the accretion disk is well converged into a steady state. This can be seen by the mass-accretion rate decelerating its decline toward the end of the simulations (see Fig. 2). For comparison, we also show $\dot{M}_{\rm{5M}}$ in the aforementioned interval in Fig. 17. The nearly identical average mass-accretion rates also aid the comparison of the efficiencies (see Eq. (18)).
We explicitly did not use the mass-accretion rate at the event horizon, $\dot{M}$ (compare Fig. 2); we instead calculated the flux at $5\,\mathrm{M}$ since this excludes the injected matter to set the density floor at the inner simulation boundary. The radial rest-mass energy density flux is defined as follows:
$\dot{M}_{\mathrm{5M}}=\int\hskip-6.0pt\int_{r=\mathrm{5M}}\sqrt{-g}\mathrm{d}\theta\mathrm{d}\phi\,\rho u^{r}.$ (19)
The resulting efficiencies of the energy-extraction process in all three models are shown in Fig. 5. On average, the K.99 simulation exhibits a vastly improved efficiency of $\sim 60\,\%$ excess compared to the commonly used MAD setup, K.80. The averages and peak values of all simulations can be found in Tab. 1.
@@
Table 1: Time averages, corresponding error, and maxima of efficiencies comparing all simulations (cf. Fig. 5).
2.4 Angular-momentum transport
In this section, we investigate the angular-momentum transport by linking it to the jet efficiency and to the impact of the magnetic energy content in the torus on the dynamics from the initialization of the simulations. Following the approach and definitions of the angular-momentum fluxes of Chatterjee and Narayan (2022), we define the radial component of the time-averaged and shell-integrated angular-momentum flux vectors, $\dot{J}^{i}$ ($i\in(r,\theta)$), accordingly. The total flux is defined as
@@ 2.4 Angular-momentum transport
and the advective flux reads
@@
The stress-induced flux is given by
@@
which can be further subdivided into the flux induced by Maxwell-stresses,
@@
and by Reynolds stresses, which are subdominant and therefore do not add anything to the discussions in this paper; hence, $\dot{J}^{i}_{\rm{stress}}\sim\dot{J}^{i}_{\rm{stress,M}}$.
Figure 6: Left panel: Time- and azimuthally averaged density distribution in the meridional plane and total angular-momentum flux streamlines (color-coded by a log-scale of the absolute value; increasing from black to white; simulation: K.80). Middle column: Polar slices of the radial angular-momentum flux at $r=2\,\mathrm{M}$, indicated by the dashed gray lines in the left panel and right column. Shown are the total, advective and Maxwell-stress-induced angular-momentum fluxes from top to bottom. Right column: Shell-integrated radial component of the angular-momentum fluxes. Solid lines indicate positive values, i.e., outward fluxes and dashed lines are the absolute values of inward (negative) flux.
@@
We time-averaged the quantities again over the $t=[25000,\,30000]\,\rm{M}$ interval in order to compare it with the jet energetics and efficiencies from Sect. 2.2 and Sect. 2.3.
@@
In Fig. 6, we present the polar and radial dependence of time- and azimuthally averaged angular-momentum transport. In the left panel, we show the time- and azimuthally averaged total angular-momentum vector field for the first time for simulations in a MAD setup with a high prograde BH spin (cf. Chatterjee and Narayan (2022) Fig. 10, showing the case of a non-rotating BH in the MAD-setup). The meridional cross-section of the angular-momentum vector field exhibits quite distinct characteristics from the case of the MAD setup with a non-rotating BH, in that it exhibits a monopole field geometry in the jet and wind region up to a polar angle of $\sim 70^{\circ}$ and transitioning into a quadrupole field geometry in the disk and torus. Through this magnetospheric configuration, the polar outflow from the jet and wind feeds back into the equatorial inflow in the disk, as opposed to the non-rotating MADs, where there is a transition from an equatorially parallel inflow inside the disk scale height (similar opening angle to the transition to the loops) to the parabolically outflowing angular-momentum flux in the jet. The major difference here is the dominance of the angular momentum transport driven by the rotating BH, where the dominantly radial outflow is redirected toward the equatorial plane into the disk, where the torus has to facilitate the radially outward angular-momentum transport in the wind region.
@@
In the middle and right columns of Fig. 6, we display the impact of the magnetic filling of the MAD torus on the angular momentum transport. The middle column shows the polar dependence of the time- and azimuthally averaged radial component of the angular-momentum vectors at $r=2$ M, where there is a maximum in the shell-integrated fluxes. The right panel shows the shell-integrated (see Sect. 2.2 for the definition of shell-integration) and time-averaged radial angular-momentum fluxes. The polar slices (indicated by dashed gray lines in the left panel and right column) of $\dot{J}_{\rm{tot}}^{r}$ and $\dot{J}_{\rm{stress,M}}^{r}$ show a clear model trend. Higher initial magnetic energy from the torus leads to an increase in outward angular-momentum flux, which is also sustained at larger radii of $r\lesssim 50\,\mathrm{M}$ (presented in Fig. 20).
@@
In the right panel at the top, one can see that the total shell-integrated fluxes are positive; i.e., radially outgoing flux in the jet and wind are dominant, whereas for the case of a non-rotating BH in Chatterjee and Narayan (2022), Fig. 14, the shell-integrated radial component of the total angular-momentum flux is constant and negative up to $r\sim 100$ M. Secondly, overall the angular-momentum fluxes peak close to the event horizon of the BH, specifically inside the ergosphere, and they drop outside of this region.
@@
We note that the model trend is not identical for the shell-integrated total and Maxwell-induced fluxes, since these are affected by the inward advective flux, which is more sensitive to turbulent realizations and does not appear to exhibit any model trend. The model trend is the clearest in the jet sheath as opposed to the equatorial plane in both total and Maxwell-induced angular-momentum fluxes. Looking at the middle panel, one can clearly see that the advective equatorial flux for K.80 is the highest; hence, it decreases the total shell-integrated flux the most for this model. However, we observe that simulation K.99 overall exhibits the highest angular-momentum flux both inward, driven by advective stresses, and especially outward, where the electromagnetic stresses are dominant. Furthermore, we observe that the model trend in the jet efficiency and the shell-integrated radial angular-momentum fluxes is consistent with one another, which directly links the efficiencies to the angular-momentum transport of these simulations.